Orwell building supplies' last dividend was $1.75. Its dividend growth rate is expected to be constant at 20.00% for 2 years, after which dividends are expected to grow at a rate of 6% forever. Its required return (rs) is 12%. What is the best estimate of the current stock price?
Answer #1

D1=(1.75*1.2)=2.1

D2=(2.1*1.2)=2.52

Value after year 2=(D2*Growth rate)/(Required return-Growth rate)

=(2.52*1.06)/(0.12-0.06)

=44.52

Hence current price=Future dividend and value*Present value of discounting factor(rate%,time period)

=2.1/1.12+2.52/1.12^2+44.52/1.12^2

=$39.375